OpenStudy (anonymous):

Three terms of a geometric sequence are shown. f(1) = 1, f(3) = 9, f(6) = –243 Which recursive formula could define the sequence? a. f(n + 1) = 9f(n) b. f(n + 1) = –9f(n) c. f(n + 1) = 3f(n) d. f(n + 1) = –3f(n) Need Help! Don't Give Answer

OpenStudy (anonymous):

@ganeshie8

OpenStudy (tkhunny):

Have you considered calculating the first 6 terms of each? a. f(n + 1) = 9f(n) f(1) = 1 f(2) = 9f(1) = 9(1) = 9 f(3) = 9f(2) = 9(9) = 81 -- Well, that's no good. We were hoping for 9. You do the rest.

OpenStudy (paxpolaris):

@justsmile531 are you set? What part do you have trouble with?

OpenStudy (anonymous):

@tkhunny Can you explain how you did that?

OpenStudy (paxpolaris):

can you tell me in your own words what a geometric sequence is?

OpenStudy (tkhunny):

I used the "a" formula definition, starting with n = 1.

OpenStudy (anonymous):

Can you explain step by step @tkhunny

OpenStudy (tkhunny):

I did already. 'a' formula definition. a. f(n + 1) = 9f(n) n = 1 f(1+1) = 9f(1) f(2) = 9f(1) We are given f(1) f(2) = 9(1) = 9 Now, we know f(2). Use this to find f(3) by substituting n = 2. etc.

OpenStudy (anonymous):

So A is not the answer; Correct ? @tkhunny

OpenStudy (anonymous):

Can you dummy it down please? @PaxPolaris

OpenStudy (paxpolaris):

a Geometric sequence is when you keep multiplying the same number (the common ratio) ... to get the next number in the sequence. correct?

OpenStudy (anonymous):

Is it C

OpenStudy (paxpolaris):

if you keep multiplying 3 .... you will never get a negative number in your sequence.... (it can't have -243)

OpenStudy (anonymous):

So D

OpenStudy (paxpolaris):

1, ___, 9, ___, ___, -243

OpenStudy (paxpolaris):

correct

OpenStudy (anonymous):

Thank you

OpenStudy (paxpolaris):

1, -3 , 9, -27, 81, -243...